5. Kilauea in Hawaii is the world’s most continuously active volcano. Very active volcanoes characteristically eject red-hot rocks and lava rather than smoke and ash. Suppose a large rock is ejected from the volcano with a speed of 25.0 m/s and at an angle 35.0º above the horizontal. The rock strikes the side of the volcano at an altitude 20.0 m lower than its starting point.

a. Calculate the time it takes the rock to follow this path.

b. What are the magnitude and direction of the rock’s velocity at impact?


1
Expert's answer
2021-11-11T10:37:34-0500

a.

"a=gsin35\degree"

"v=gtsin35\degree+v_0"

"d=gt^2sin35\degree/2+v_ot=20/sin35\degree"


"2.8t^2+25t-34.9=0"


"t=\frac{-25+\sqrt{625+4\cdot2.8\cdot34.9}}{5.6}=1.23" s


b.

"v=9.8\cdot1.23sin35\degree+25=31.9" m/s

direction: 35.0º above the horizontal